# The emergence of inherently 9-dimensional one-loop couplings from T-duality

### v1

@@
one-loop couplings from T-duality
@@
Recent studies suggest that applying the Buscher rules to the dimensional reduction of ten-dimensional, one-loop effective actions may generate ”purely stringy” couplings in nine dimensions that cannot be lifted to a local, covariant form in ten dimensions. We investigate this phenomenon at order $\alpha^{\prime 3}$ in type IIA string theory. By computing the circular reduction of the one-loop Chern-Simons term and pure-gravity couplings in type IIA theory and applying the T-duality transformation to the resulting couplings, we derive their counterparts in the type IIB effective action. We demonstrate that the resulting nine-dimensional type IIB couplings are invariant under S-duality without requiring contributions from the tree-level effective action or non-perturbative effects. This result is non-trivial: while S-duality forbids one-loop Chern-Simons couplings in ten-dimensional type IIB theory, we show that the one-loop Chern-Simons couplings in nine dimensions, derived from the T-duality of type IIA, are S-duality invariant. As a consistency check, we show that the nine-dimensional type IIB couplings, when reduced on a K3 surface, reproduce the known heterotic string couplings on $T^{5}$ at order $\alpha^{\prime}$, via the duality between the two theories.
@@ 1 Introduction
The spacetime effective action in string theory at the critical dimension 10 is characterized by a double expansion in the world-sheet genus $g$ and the string scale $\alpha^{\prime}$, which governs the derivative expansion. A range of complementary techniques—including the S-matrix method [1, 2, 3], the non-linear sigma model [4, 5, 6], supersymmetry [7], T-duality [8, 9, 10], and S-duality [11, 12]—are employed to determine the coefficients of this expansion. The dualities impose distinct constraints: T-duality relates classical effective actions at different orders in $\alpha^{\prime}$ while preserving the genus, whereas S-duality operates more profoundly, connecting 10-dimensional couplings across different genera and introducing non-perturbative effects. A striking application of this principle is that imposing S-duality on the classical type IIB effective action at order $\alpha^{\prime 3}$ uniquely determines the structure of its higher-genus and non-perturbative completions [12].
@@
To impose the T-duality constraint on the effective action, it is necessary to compactify at least one spatial dimension on a circle. If the circle’s radius is treated as a dimensionless parameter, the resulting nine-dimensional theory acquires an additional expansion parameter alongside the $\alpha^{\prime}$- and $g$-expansions: an expansion in the circle’s radius [13]. At the classical level, the effective action is background-independent [14]. The nine-dimensional effective action can therefore be derived from the Kaluza-Klein (KK) reduction of the ten-dimensional theory. Consequently, the radius-dependence of the classical couplings is known exactly from this KK reduction. This property, in fact, allows one to determine the classical effective action by imposing T-duality on its KK reduction [15, 16, 17, 18]. At the loop level, however, the effective action is background-dependent [13]. The nine-dimensional effective action is then distinctly different from its ten-dimensional counterpart, and the radius-dependence of the couplings generally cannot be derived from the KK reduction of the ten-dimensional terms. In this case, the couplings must be expressed as a radius expansion. The leading large-radius term of this expansion can be found from the KK reduction of the ten-dimensional couplings. Wherease, the leading small-radius terms can be found by applying the T-duality transformation to the leading large-radius term . These small-radius terms, along with all other possible terms in the radius expansion, represent inherently nine-dimensional couplings that cannot be derived from the KK reduction of the ten-dimensional theory [13]. They may be determined through explicit S-matrix calculations that include the effects of compactification on a circle [19]. Ultimately, the radius expansion must be constructed to ensure the loop-level effective action is invariant under T-duality.
@@
This paper derives the nine-dimensional, one-loop effective action of type IIB theory at order $\alpha^{\prime 3}$ in the small-radius limit by imposing T-duality on the circularly reduced ten-dimensional action of type IIA theory. We demonstrate that, while the large-radius limit of the type IIB Chern-Simons term is zero, its small-radius limit is non-zero and is, in fact, invariant under S-duality. Furthermore, whereas the large-radius limit of the type IIB one-loop pure gravity effective action must combine with tree-level and non-perturbative couplings to achieve S-duality invariance [12], we show that its leading small-radius couplings are S-duality invariant by themselves. We establish this by explicitly reducing the ten-dimensional couplings from type IIA and performing a T-duality transformation to derive the corresponding nine-dimensional couplings in type IIB. We then show that the resulting couplings are invariant under S-duality when the Ramond-Ramond (RR) fields are set to zero.
@@
We observe that S-duality constrains the radius expansion of the type II effective action at order $\alpha^{\prime 3}$, predicting its truncation to solely the large- and small-radius limits. In the nine-dimensional type IIB theory, the one-loop couplings at this order exhibit S-duality in specific regimes: in the large-radius limit, they combine with tree-level and non-perturbative contributions to achieve invariance [12], while in the small-radius limit, we demonstrate that they are inherently invariant. Given that the tree-level action and non-perturbative effects are already incorporated, S-duality requires that any remaining terms in the one-loop radius expansion must themselves be invariant. The only consistent solution to this constraint is for such terms to vanish, thereby confirming the truncation of the radius expansion.
@@
As a further check on the nine-dimensional type IIB couplings at order $\alpha^{\prime 3}$, we use the duality between type IIB theory on $S^{1}\times K3$ and heterotic theory on $T^{5}$ [20]. We dimensionally reduce the couplings on a K3 manifold and demonstrate that, while their large-radius limit yields no terms at order $\alpha^{\prime}$, their small-radius limit produces five-dimensional couplings at order $\alpha^{\prime}$. These are shown to be consistent with the expected couplings of the heterotic theory compactified on $T^{5}$, following an appropriate field redefinition that maps the type IIB radius parameter to the heterotic dilaton. Given that the dilaton dependence in the heterotic theory at order $\alpha^{\prime}$—specifically, the factor of $e^{-2\Phi}$—is exact [21, 22], this correspondence implies that the radius dependence of the type IIB couplings is also exact. This result reinforces the conclusion that the nine-dimensional couplings exist in only two distinct forms: one for the large-radius limit and one for the small-radius limit, with no interpolating corrections.
@@
This paper is structured as follows. Section 2 details the KK reduction of the type IIA Chern-Simons term at order $\alpha^{\prime 3}$, its transformation under T-duality to produce the appropriate couplings in type IIB, and the invariance of the resulting couplings under S-duality. In subsection 2.4, we show that S-duality prohibits the existence of Chern-Simons couplings in ten dimensions. In subsection 2.5, we demonstrate that the K3 reduction of the nine-dimensional couplings in type IIB produces four-derivative couplings consistent with the corresponding heterotic theory on $T^{5}$. In Section 3, we perform the same calculations for the KK reduction of the pure gravity couplings in type IIA and study their transformation under T-duality and S-duality, as well as the equivalence of their K3 reduction of nine-dimensional type IIB couplings with the heterotic theory on $T^{5}$. Our conclusions are presented in Section 4. All calculations were performed using the “xAct” package [23].
@@ 2 Chern-Simons term at order $\alpha^{\prime 3}$
The KK reduction of the eight-derivative Chern-Simons term in M-theory produces the corresponding Chern-Simons term in type IIA theory [27], as well as some gauge-invariant couplings involving RR fields [33]. The latter couplings are not discussed here, as our focus is on NS-NS couplings. The Chern-Simons term in type IIA theory is (see [33] for our notation and conventions)
@@
where $\epsilon_{10}$ denotes the Levi-Civita tensor in ten dimensions, and $\kappa^{2}=\frac{1}{\pi}(2\pi\sqrt{\alpha^{\prime}})^{8}$. The absence of a dilaton in the above action identifies it as the one-loop effective action. Since 11-dimensional M-theory lacks a fundamental string, the KK reduction of its effective action correctly produces the 10-dimensional type IIA couplings. In other words, the above action is valid for any radius of the 11th direction $R_{11}=g_{s}\sqrt{\alpha^{\prime}}$. However, this is not the case for the KK reduction of type IIA theory, whose fundamental object is a string originating from an M2-brane wrapped on the circle of the 11th dimension. Consequently, the lower-dimensional one-loop effective action of type IIA string theory may receive both winding and KK contributions. The KK contribution, which we calculate in the following section, is valid only in the large-radius limit.
@@
where indices $a,b$ denote directions orthogonal to the Killing coordinate $y$. The resulting nine-dimensional base space fields are: $\bar{g}_{ab}$ (metric), $\bar{b}_{ab}$ (antisymmetric tensor), $\bar{\phi}$ (dilaton), $\varphi$ (scalar), and the vector fields $g_{a}$ and $b_{a}$.
@@
The action in equation (4), which represents the one-loop effective action of type IIA theory only in the large-radius limit, can be rederived by studying the appropriate torus-level S-matrix elements. When calculating S-matrix elements in type IIA string theory compactified on a circle, the continuous spacetime momentum in the circle direction is replaced with discrete KK and winding momenta [19]. Consequently, the corresponding spacetime couplings may, in general, receive contributions from both winding and KK modes. However, the couplings in (4) are reproduced solely by the pure KK modes. It has been speculated in [13] that because $S^{\rm CS}_{\rm IIA}$ is topological, there is no overall radius factor in this action; hence, the nine-dimensional S-matrix calculation should not produce any contact terms that involve winding modes in type IIA theory. In other words, the radius dependence of the couplings in (4) is exact. This implies there is no small-radius limit or winding-mode contribution for the type IIA couplings in (4).
Although there are no couplings corresponding to winding modes in the 9-dimensional type IIA Chern-Simons terms (4), such couplings do exist in the 9-dimensional type IIB theory, where they correspond to pure winding modes in its S-matrix elements. In this case, one expects there to be no couplings corresponding to KK modes instead. The couplings for the pure winding contribution in type IIB theory can be derived via T-duality from the type IIA couplings, which we will study in the next section.
@@ 2.3.1 Consistency with S-duality
and the Lagrangian density $\mathcal{L}_{bcdefghi}$ contains four Riemann tensors and four inverse metrics.
@@
Therefore, the type IIB couplings given in (10) are invariant under the continuous $SL(2,\mathbb{R})$ S-duality group, just like the two-derivative effective action of type IIB (see e.g., [25]). That is,
@@
It has been speculated in [13] that because $S_{\rm IIB}^{CS}$ is a topological coupling, its radius dependence is exact; in other words, the action contains no overall function of the radius. This implies that for the corresponding S-matrix elements in type IIB string theory—which involve both KK and winding momentum in the circular direction—the terms with KK momentum produce no contact terms. The only contact terms arise in the limit involving purely winding momentum. Conversely, if the action (10) contained an overall function of the radius, it would not be invariant under S-duality. Hence, consistency with S-duality confirms that the radius dependence in the nine-dimensional couplings (10) is exact. Moreover, the observation that the type IIB couplings (10) correspond solely to winding momentum in the S-matrix elements predicts that S-matrix elements in 10-dimensional Minkowski spacetime—which have no winding momentum—should produce no contact terms. This implies that there can be no Chern-Simons term in the 10-dimensional type IIB theory. In the following section, we confirm this prediction by studying S-duality in 10-dimensional spacetime.
2.4 No 10D Chern-Simons term in type IIB
It has been observed that the Chern-Simons coupling in type IIA theory, which involves an odd number of B-fields, is absent in type IIB theory [30]. The above discussion indicates that while there exist nine-dimensional Chern-Simons terms involving an even number of $W$ and $\bar{H}$—which are reductions of the $B$-field strength—there should be no ten-dimensional Chern-Simons term in type IIB theory involving an even number of B-fields. In fact, such couplings would be inconsistent with S-duality.
First, note that a Chern-Simons term cannot be constructed with zero $B$-field. To show that no such term exists with two $B$-fields, consider a potential Chern-Simons coupling with two $H$-fields in the string frame:
$\epsilon^{\prime}_{10}HHRRR$ (26)
where $\epsilon^{\prime}_{10}$ is the 10-dimensional Levi-Civita symbol. The above coupling has three Riemann curvatures and four inverse metrics. Hence, it appears in the Einstein frame as
$\epsilon^{\prime}_{10}e^{-\Phi/2}HHR^{E}R^{E}R^{E}.$ (27)
Using the relation (16), one finds that it is not invariant under S-duality. Similarly, consider the following coupling in the string frame:
$\epsilon^{\prime}_{10}\nabla H\nabla HRR$ (28)
It has two Riemann curvatures and three inverse metrics; hence, in the Einstein frame, it appears as
$\epsilon^{\prime}_{10}e^{-\Phi/2}\nabla H\nabla HR^{E}R^{E}$ (29)
which is not invariant under S-duality. Similar arguments hold for other terms involving four, six, and eight H-fields. For the theory to be $SL(2,\mathbb{Z})$-invariant, any one-loop coupling would need to be combined with a corresponding tree-level Chern-Simons term, along with non-perturbative corrections. The fact that no such tree-level term exists implies that S-duality forbids a 10-dimensional Chern-Simons term in type IIB string theory, which is consistent with expectations.
@@ 3 Pure gravity couplings at order $\alpha^{\prime 3}$
In the next subsection, we find the KK reduction of the above coupling. Since the above action is the one-loop effective action of type IIA string theory, the KK reduction does not in general produce the complete couplings in nine dimensions. There may be winding contributions as well. Unlike the Chern-Simons term in type IIA theory, which has no contributions from winding modes, the pure gravity part does have winding contributions. These can be found by applying T-duality to the KK reduction of the one-loop effective action of type IIB theory, which is given by
@@ 3.1 Large-radius limit of 9D couplings in type IIA
The coefficients in this work are determined using the scheme defined in the Appendix (see (4)), where (3.1) is equated to our minimal basis after applying field redefinitions, integration by parts, and Bianchi identities. That is,
@@
where the Lagrangian $\mathcal{L}(\varphi,V)$ is given in (4), with the coupling constants set to the values in (42). Since this action is derived from the KK reduction of the ten-dimensional couplings in (3), it is valid only in the large-radius limit. The corresponding couplings in the small-radius limit can be found via T-duality from the KK reduction of the type IIB couplings in (39), which is not the focus of this work. In the next section, we instead derive the small-radius limit of the type IIB couplings at order $\alpha^{\prime 3}$.

### v2

# The emergence of inherently 9-dimensional one-loop effective action from T-duality

@@
one-loop effective action from T-duality
@@
Recent studies suggest that applying the Buscher rules to the dimensional reduction of ten-dimensional, one-loop effective actions generate ”purely stringy” couplings in nine dimensions that cannot be lifted to a local, covariant form in ten dimensions. We investigate this phenomenon at order $\alpha^{\prime 3}$ in type IIA string theory. By computing the circular reduction of the one-loop Chern-Simons term and pure-gravity couplings in type IIA theory and applying the T-duality transformation to the resulting couplings, we derive their counterparts in the type IIB effective action. We demonstrate that the resulting nine-dimensional type IIB couplings are invariant under S-duality without requiring contributions from the tree-level effective action or non-perturbative effects. As a consistency check, we show that the nine-dimensional type IIB couplings, when reduced on a K3 surface, reproduce the known heterotic string couplings on $T^{5}$ at order $\alpha^{\prime}$, via the duality between the two theories.
@@ 1 Introduction
The spacetime effective action in string theory at the critical dimension 10 is characterized by a double expansion in the world-sheet genus $g$ and the string scale $\alpha^{\prime}$, which governs the derivative expansion. A range of complementary techniques—including the S-matrix method [1, 2, 3], the non-linear sigma model [4, 5, 6], supersymmetry [7], T-duality [8, 9, 10], and S-duality [11, 12]—are employed to determine the coefficients of this expansion. The dualities impose distinct constraints: T-duality relates effective actions at different orders in $\alpha^{\prime}$ while preserving the genus, whereas S-duality operates more profoundly, connecting 10-dimensional couplings across different genera and introducing non-perturbative effects. A striking application of this principle is that imposing S-duality on the classical type IIB effective action at order $\alpha^{\prime 3}$ uniquely determines the structure of its higher-genus and non-perturbative completions [12, 13].
@@
To impose the T-duality constraint on the effective action, it is necessary to compactify at least one spatial dimension on a circle. If the circle’s radius is treated as a dimensionless parameter, the resulting nine-dimensional theory acquires an additional expansion parameter alongside the $\alpha^{\prime}$- and $g$-expansions: an expansion in the circle’s radius [14]. At the classical level, the effective action is background-independent [15]. The nine-dimensional effective action can therefore be derived from the Kaluza-Klein (KK) reduction of the ten-dimensional theory. Consequently, the radius-dependence of the classical couplings is known exactly from this KK reduction. This property, in fact, allows one to determine the classical effective action by imposing T-duality on its KK reduction [16, 17, 18, 19]. At the loop level, however, the effective action is background-dependent [14]. The nine-dimensional effective action is then distinctly different from its ten-dimensional counterpart, and the radius-dependence of the couplings generally cannot be derived from the KK reduction of the ten-dimensional terms. In this case, the couplings must be expressed as a radius expansion.
The leading large-radius term of this expansion is obtained from the KK reduction of the ten-dimensional couplings. In contrast, the leading small-radius terms follow by applying a T-duality transformation to this large-radius result. These small-radius terms—along with all other possible terms in the radius expansion—represent inherently nine-dimensional couplings that cannot be derived from the KK reduction of the ten-dimensional theory [14]. They may instead be determined through explicit S-matrix calculations incorporating the effects of compactification on a circle. This compactification replaces continuous momentum in one direction with discrete KK momentum, while also introducing winding modes along the circle [20]. While the KK momentum contribution to the effective action originates from the KK reduction of ten-dimensional couplings, the winding mode contribution cannot be found this way. Ultimately, the full radius expansion must be constructed to ensure the nine-dimensional, loop-level effective action is invariant under T-duality [14].
@@
By applying T-duality to the circularly reduced, ten-dimensional one-loop effective action of type IIA theory, we derive the corresponding nine-dimensional effective action for type IIB theory at order $\alpha^{\prime 3}$ in the small-radius limit. We show that, whereas S-duality in the large-radius limit requires the inclusion of tree-level and non-perturbative effects [12, 13, 21], the leading small-radius couplings we find exhibit S-duality invariance independently. This invariance is explicitly verified by performing the dimensional reduction and T-duality transformation, and subsequently demonstrating that the resulting nine-dimensional type IIB couplings are S-duality invariant in the absence of Ramond-Ramond (RR) fields.
@@
A key observation is that S-duality forces the radius expansion of the type II effective action at order $\alpha^{\prime 3}$ to truncate, leaving only the large- and small-radius limits [14]. The argument proceeds from the known S-duality properties of the nine-dimensional type IIB one-loop couplings at this order. While these couplings require combination with tree-level and non-perturbative terms to be invariant in the large-radius limit [12, 13, 21], we demonstrate they are inherently invariant in the small-radius limit. Given that the tree-level and non-perturbative sectors are already included, any further one-loop terms in a full radius expansion would, by S-duality, need to be separately invariant. The sole consistent possibility is for such terms to be absent, which confirms the predicted truncation.
@@
As a further check on the nine-dimensional type IIB couplings at order $\alpha^{\prime 3}$, we use the duality between type IIB theory on $S^{1}\times K3$ and heterotic theory on $T^{5}$ [22, 23]. We dimensionally reduce the couplings on a K3 manifold and demonstrate that, while their large-radius limit yields no terms at order $\alpha^{\prime}$, their small-radius limit produces five-dimensional couplings at order $\alpha^{\prime}$. These are shown to be consistent with the expected couplings of the heterotic theory compactified on $T^{5}$, following an appropriate field redefinition that maps the type IIB radius parameter to the heterotic dilaton. Given that the dilaton dependence in the heterotic theory at order $\alpha^{\prime}$—specifically, the factor of $e^{-2\Phi}$—is exact [24, 25], this correspondence implies that the radius dependence of the type IIB couplings is also exact. This result reinforces the conclusion that the nine-dimensional couplings exist in only two distinct forms: one for the large-radius limit and one for the small-radius limit, with no interpolating corrections.
@@
This paper is structured as follows. In Section 2, we detail the KK reduction of the type IIA Chern-Simons term at order $\alpha^{\prime 3}$, perform its T-duality transformation to derive the corresponding type IIB couplings, and demonstrate the S-duality invariance of the result. Subsection 2.5 further shows that the K3 reduction of these nine-dimensional, small-radius type IIB couplings yields four-derivative couplings consistent with the heterotic theory on $T^{5}$. Section 3 extends this analysis to the pure gravity couplings of type IIA, examining their KK reduction, transformation under T-duality and S-duality, and the equivalence of their K3-reduced type IIB form with heterotic theory on $T^{5}$. Our conclusions are presented in Section 4. All computations were performed using the “xAct” package [26].
@@ 2 Chern-Simons term at order $\alpha^{\prime 3}$
The KK reduction of the eight-derivative Chern-Simons term in M-theory produces the corresponding Chern-Simons term in type IIA theory [30], as well as some gauge-invariant couplings involving RR fields [35]. The latter couplings are not discussed here, as our focus is on NS-NS couplings. The Chern-Simons term in type IIA theory is
@@
where $\epsilon_{10}$ denotes the Levi-Civita tensor in ten dimensions, and $\kappa^{2}=\frac{1}{\pi}(2\pi\sqrt{\alpha^{\prime}})^{8}$. The absence of a dilaton in the above action identifies it as the one-loop effective action. Since 11-dimensional M-theory lacks a fundamental string, the KK reduction of its effective action correctly produces the 10-dimensional type IIA couplings. In other words, the above action is valid for any radius of the 11th direction $R_{11}=g_{s}\sqrt{\alpha^{\prime}}$. However, this is not the case for the KK reduction of type IIA theory, whose fundamental object is a string originating from an M2-brane wrapped on the circle of the 11th dimension. Consequently, the lower-dimensional one-loop effective action of type IIA string theory may receive both winding and $K$ contributions. The KK contribution, which we calculate in the following section, is valid only in the large-radius limit. [^2] The work in [36] examined the KK reduction incorporating the B-field via a connection with torsion and analyzed its T-duality transformations in the zero-radion case.
@@
where indices $a,b$ denote directions orthogonal to the Killing coordinate $y$. The resulting nine-dimensional base space fields are: $\bar{g}_{ab}$ (metric), $\bar{b}_{ab}$ (antisymmetric tensor), $\bar{\phi}$ (dilaton), $\varphi$ (radion), and the vector fields $g_{a}$ and $b_{a}$.
@@
The action in equation (4), which captures the one-loop effective action of type IIA theory in the large-radius limit, can also be derived from the relevant torus-level S-matrix elements. In type IIA string theory compactified on a circle, the continuous momentum along the circle is replaced by discrete KK and winding momenta [20]. Consequently, the corresponding spacetime couplings generally receive contributions from both types of modes. The couplings in (4), however, arise exclusively from the pure KK sector. In the next section, we will use T-duality to derive from these type IIA couplings the corresponding pure winding contributions for type IIB theory.
@@ 2.3.1 Consistency with S-duality
and the Lagrangian density $\mathcal{L}_{bcdefghi}$ contains four Riemann tensors and four inverse metrics. We also confirmed that the dilaton derivative does not appear when transforming the first term in (10) to the Einstein frame.
@@
Therefore, the type IIB couplings given in (10), which are valid in the small-radius limit, are invariant under the continuous $SL(2,\mathbb{R})$ S-duality group—much like the two-derivative effective action of type IIB (see, e.g., [28]). That is,
@@
In contrast, a non-zero ten-dimensional Chern–Simons term involving an even number of $B$-fields has been proposed in [36] and shown to be S-duality invariant after including tree-level couplings with RR fields as well as non-perturbative contributions [21]. The KK reduction of these terms yields S-duality invariant couplings in the large-radius limit. Consequently, while the small-radius couplings are invariant under the continuous $SL(2,\mathbb{R})$ group, the large-radius couplings become invariant under the discrete $SL(2,\mathbb{Z})$ group only after the inclusion of tree-level and non-perturbative corrections. Because the tree-level and non-perturbative sectors are incorporated independently at large radius, S-duality forces any additional one-loop terms to be separately invariant. This condition, together with the transformation of the radion field $\varphi$ from the string frame to the Einstein frame (see (17)), leads to the conclusion that only the small-radius coupling fulfills the required invariance. The same conclusion can be reached by studying the duality in 5D that we consider in the next section.
@@ 3 Pure gravity couplings at order $\alpha^{\prime 3}$
In the next subsection, we find the KK reduction of the above coupling. Since the above action is the one-loop effective action of type IIA string theory, the KK reduction does not produce the complete couplings in nine dimensions. There should be winding contributions as well. These can be found by applying T-duality to the KK reduction of the one-loop effective action of type IIB theory, which is given by
@@ 3.1 Large-radius limit of 9D couplings in type IIA
We determine the coefficients in this work by employing the scheme outlined in the Appendix (see (4)). This involves equating the result in (3.1) to our chosen minimal basis, once field redefinitions, integration by parts, and Bianchi identities have been applied. That is,
@@
where the Lagrangian $\mathcal{L}(\varphi,V)$ is given in (4), with the coupling constants set to the values in (38). Since this action is derived from the KK reduction of the ten-dimensional couplings in (3), it is valid only in the large-radius limit. The corresponding couplings in type IIA in the small-radius limit can be found via T-duality from the KK reduction of the type IIB couplings in (35), which is not the focus of this work. In the next section, we instead derive the small-radius limit of the type IIB couplings at order $\alpha^{\prime 3}$.